ID |
Date |
Author |
Subject |
Text |
 |
299
|
Sun Feb 27 08:00:23 2022 |
RSS TD | Current status | Ring current ~7.5e7 particles/s
Merger, Tape server (WS3):
1414 kB/s writing ok |
  |
298
|
Sun Feb 27 06:58:51 2022 |
RSS TD | Current status | 7:59: 95K-->120K
8:06: 120K --> 95K
Ring current ~7e7 particles/s |
   |
297
|
Sun Feb 27 06:23:42 2022 |
RSS TD | Current status | Ring current ~6e7 particles/s
Merger, Tape server (WS3):
1414 kB/s writing ok |
6x |
296
|
Sun Feb 27 05:45:53 2022 |
David, Pierre-Michell | Current status | 6:05: 95K-->120K
6:12: 120K --> 95K
Ring current 8e7 particles/s |
   |
295
|
Sun Feb 27 04:54:46 2022 |
David, Pierre-Michell | Current status | Ring current 8e7 particles/s
Merger, Tape server (WS3):
1664 kB/s writing ok |
  |
294
|
Sun Feb 27 03:49:36 2022 |
David, Pierre-Michell | Current status | 04:06 85K -->120K : 04:13 120K -->95K
current ring 8e7 p/s
Merger, Tape server (WS3): |
  |
293
|
Sun Feb 27 02:38:52 2022 |
David, Pierre-Michell | Current status | We have decreased the temperatur> 95K
--> 90K. At 95K the target was stable,
but let's try to increase the density |
   |
292
|
Sun Feb 27 01:46:56 2022 |
David, Pierre-Michell | Current status | So far looks like with a temperature of
95K the target is maintained with a reasonable
density throughout long periods of time. |
  |
291
|
Sun Feb 27 00:53:24 2022 |
David, Pierre-Michell | Current status | We have recovered the beam at 01:49, after
such a long time we had no target so we increased
from 80K to 120K at 01:52. 02:01 update 120K-->95K |
7x |
290
|
Sun Feb 27 00:07:15 2022 |
David, Pierre-Michell | No beam, valve closed | At 00:37 we lost the beam due to a high
pressure burst in the tarjet which resulted
in a valve closed at YR09VV... |
|
289
|
Sat Feb 26 23:20:13 2022 |
David, Pierre-Michell | Update status | In this last hour we have been playing
with the cryostat temperature as dictated
by Nikos, in order to have the target working |
7x |
288
|
Sat Feb 26 22:22:27 2022 |
David, Pierre-Michell | Starting to take data | At 23:22 we started recording data under
the new conditions.
Merger, Tape server (WS3): |
11x |
287
|
Sat Feb 26 04:36:38 2022 |
RSS TD | Saturday 26 February 07:00-15:00 | 05:36 E4 pressure c. 1.7e-9mbar *not* 3e-10mbar
as reported earlier
See attachments 1 & 2
|
27x |
286
|
Fri Feb 25 19:20:22 2022 |
MS, PJW | check | 'Check Clock Status', 'Check
ADC Calibration' and 'Check the White
rabbit decoder status' no changes |
11x |
285
|
Fri Feb 25 17:48:22 2022 |
MS, PJW | check | 'Check Clock Status', 'Check
ADC Calibration' and 'Check the White
rabbit decoder status' - no changes |
11x |
284
|
Fri Feb 25 16:01:04 2022 |
MS, PJW | check |
'Check Clock Status': Clock
status test result: Passed 5, Failed 0 |
7x |
283
|
Fri Feb 25 15:38:40 2022 |
JM, CB, PJW | Run 55, 56, 59 root histograms | Run55 - detector out, start 22/2/22 13:20,
end 23/2/22 10:49
Run 56 - detector moved in, start |
6x |
282
|
Fri Feb 25 11:20:30 2022 |
CB | Pressure and temperature data saved | Saved pressure reading and temperature
data logs from internal target in /TapeServer/22Feb/OtherData
Leakage current data is saved in |
|
281
|
Fri Feb 25 09:58:08 2022 |
GHC, CB, ML, EM | Current status | 10:58
Started new run: 77
|
|
280
|
Fri Feb 25 08:14:47 2022 |
George, Esther | Current status | 9:14
Beam intensity around ~7.5e7
Pressures OK |
|